Let f(x)=2x+2 and g(x)=2x^2+4x.
After simplifying, (f+g)(x)=
solve: (f+g)(x)=

Answer by Theo(13342)

f(x) = 2x+2

g(x) = 2x^2 + 4x

f+g(x) = f(x) + g(x) = 2x + 2 + 2x^2 + 4x

combine like terms on the right side of the equation to get:

f+g(x) = f(x) + g(x) = 6x + 2 + 2x^2

reorder the terms in descending order of degree to get:

f+g(x) = f(x) + g(x) = 2x^2 + 6x + 2